A ball is thrown straight upward with a speed, v from a point, h meters above the ground. Show that the time taken, t for the ball to strike the ground is t= (v/g)[1+√1+(2gh/v^2)]


1
Expert's answer
2021-02-09T13:34:36-0500
"h(t)=h_o+v_0t-\frac {gt^2}{2}"

"h(t)=0"


"gt^2-2vt-2h=0"


"t=\frac {v+\sqrt {v^2+2gh}}{g}=\frac {v}{g}+\sqrt{\frac {v^2+2gh}{g^2}}"